# Annotation Free Semantic Segmentation with Vision Foundation Models

### v1

@@ 4.2 Annotation Free Semantic Segmentation Results
Table 1 reports the mIoU on the different datasets. Results of methods in first block are taken from TCL [7] while second block methods are taken from their corresponding papers.
@@
First, training free methods, CLIP-DIY [33], SCCLIP [31], CaR [30] and ReCo [29] perform inferior to our method mostly with a large margin except for CaR [30] on PV-21 which can be attributed to the post-processing of background in CaR. Second, FMbSeg performs best or second best on all datasets even improving over TCL [7] trained on 15M images with image level labels. FMbSeg outperforms TCL by a large margin of 8% on PC-59 and 26% on CO-80. Third, FMbSeg outperforms SAM-CLIP [32] significantly on PV-21 and PC-59 while being inferior on CS-171 and ADE albeit with a small margin. SAM-CLIP [32] trains SAM encoder on 41M images with no code or model available which makes it infeasible for us to evaluate it on the remaining datasets. Nevertheless, we emphasize that our alignment module is light and trained only for a few epochs on almost 400 times less data compared to  [32]. This makes our method a plug and play approach to semantically segment any dataset.
@@
Due to space limits in the main paper, we extensively ablate different ways of leveraging our approach in Appendix. We first consider the stage.1 only 3.2 as a training free method,  FMbSeg Stage.1.1 and  FMbSeg Stage.1.2. We also consider a semi supervised variant where image level labels are provided to the first stage. Stage.1.2 alone results in inferior results to Stage.1.1, however, the results validate that Stage.1.2 helps avoiding some inherit mistakes of Stage.1.1, leading to a boost in performance when combined. Besides, image-level labels can significantly improve the Stage 1 performance leading to SOTA performance without any training. We further ( in Appendix) examine our alignment module (pretrained on COCO-Stuff) when fine-tuned on specific target datasets mentioned in table 1 with no ground-truth annotations and in a self-supervised manner. This fine-tuning does not add a large over-head due to the efficiency of our lightweight alignment module. As shown in Appendix, in this case, the performance systematically exceeds all self-supervised methods on all datasets.
@@ 4.3 Qualitative Results
First we inspect the alignment of patch features with vocabulary after we train our alignment module. Fig. 3 shows the patch-level similarity between the image and the image-level label on Pascal VOC. Notice that the newly aligned image encoder produces more consistent similarity scores for the given categories than those produced by CLIP. This shows the efficacy of aligning a better vision encoder with text rather than trying to improve CLIP patches features which typically requires large scale (re)training.
Next, Fig. 4 shows qualitative results of the annotation free segmentation by FMbSeg on the different datasets.
4.4 Ablation
4.4.1 Choice of Loss Function
In the method section 3.3 we propose a loss function, adapted from [2], that applies contrastive learning while treating text features as prototypes. Here we ablate the effectiveness of the two terms: 1) The supervised contrastive loss [18] (SupCon) alone where text features are considered as examples of the corresponding concepts similar to the image patches of a given category. 2) The prototype loss alone Eq (2). Table 2 reports the mIOU on CS-171. We find that SupCon term alone exhibits poor convergence, with long epochs needed, leading to sub-optimal performance. The prototype loss shows a a better performance and faster convergence, probably due to the smooth approximation of Cross Entropy loss and SupCon loss on text-patch pairs. However, when combined with SupCon on pairs of patches, better performance is shown due to more enhancement in the expressiveness of these patches features. Overall, the unique treatment of TSupCon to the text features allows a smoother generalization, better convergence and hence considerably better performance.
4.4.2 Choice of Architecture
We design our alignment module as a single transformer block where multi-head self-attention is applied over image patches. We ablate our choice against other designs, namely a single linear layer and a Multi-layer perceptron (MLP) with GELU activation. Table 2 reports reports the mIOU on CS-171. The differences are not substantial, with MLP achieving better better performance than linear. The transformer block further improves over the MLP.
Module Arch. CS-171
mIoU
Linear 25.32
MLP 26.43
Transformer block 27.61
Alignment Loss CS-171
mIoU
$\ell_{SupCon}$ 24.61
[…]
$\ell_{\text{TSupCon}}$(1) 27.61
Table 2: Left. Comparison of different design choices for our alignment module, a transformer block has a small advantage. Right. Comparison of different losses. SupCon alone is inferior to our full loss TSupCon performing the best.
@@ 5 Conclusion and Future Work
Given the impressive performance of vision-language foundation models and their transferability to various downstream tasks, we consider the problem of open vocabulary annotation free semantic segmentation. By composition of different foundation models, namely CLIP and SAM, we extract free pixel-level annotations. We then propose a lightweight alignment module that projects the embedding of any arbitrary pretrained vision encoder to the text encoder space. By pretraining on COCO-Stuff, we obtain a generalist alignment module that can perform smantic segmentation across a wide range of object categories. Alternatively, our method can be deployed as a plug-and-play customized alignment module for any semantic segmentation dataset with zero annotations. We show SOTA results and open a new direction for zero-shot semantic segmentation by foundational models compositionality. As future work, we want to investigate other image encoders and continuous finetuning on new categories. Additionally, we want to explore how our alignment module can improve VLM models that are based on CLIP to further strengthen their object localization capabilities.

### v2

@@ 4.2 Annotation-Free Semantic Segmentation Results
Table 1 reports the mIoU on the different datasets. Results of methods in first block are taken from TCL [Cha et al.(2023)Cha, Mun, and Roh] while second block methods are taken from their corresponding papers.
@@
First, training free methods, CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni], SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille], CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] and ReCo [Shin et al.(2022)Shin, Xie, and Albanie] perform inferior to our method mostly with a large margin except for CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] on PV-21 which can be attributed to the post-processing of background in CaR. Second, FMbSeg performs best or second best on all datasets even improving over TCL [Cha et al.(2023)Cha, Mun, and Roh] trained on 15M images with image level labels. FMbSeg outperforms TCL by a large margin of 8% on PC-59 and 26% on CO-80. Third, FMbSeg outperforms SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] significantly on PV-21 and PC-59 while being inferior on CS-171 and ADE albeit with a small margin. SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] trains SAM encoder on 41M images with no code or model available which makes it infeasible for us to evaluate it on the remaining datasets. Nevertheless, we emphasize that our alignment module is light and trained only for a few epochs on almost 400 times less data compared to  [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari]. This makes our method a plug and play approach to semantically segment any dataset.
@@
Due to space limits in the main paper, we extensively ablate different ways of leveraging our approach in Appendix. We evaluate stage 1.1 and 1.2 performance as training free methods. We also consider a semi supervised variant where image level labels are provided to stage 1.1 which significantly improves the performance leading to SOTA performance without any training. In Appendix we further examine our alignment module (pretrained on COCO-Stuff) when fine-tuned on specific target datasets mentioned in table 1 with no ground-truth annotations and in a self-supervised manner. This fine-tuning does not add a large over-head due to the efficiency of our lightweight alignment module. As shown in Appendix, in this case, the performance on ADE20k increases to 21.57%.
@@ 4.3 Qualitative Results
First we inspect the alignment of patch features with vocabulary after we train our alignment module. Fig. 2 shows the patch level similarity between the image features and the image level text features on Pascal VOC. Notice that the newly aligned image encoder produces distinctly more consistent similarity heatmaps for the given categories than those produced by CLIP. This shows the efficacy of aligning a better vision encoder with text rather than trying to improve CLIP patches features which typically requires large scale (re)training.
Next, Fig. 4 shows qualitative results of the annotation-free zeroshot segmentation by FMbSeg on the different datasets.
@@ 5 Conclusion and Future Work
Given the impressive performance of vision-language foundation models and their transferability to various downstream tasks, we consider the problem of open vocabulary annotation-free semantic segmentation. By composition of different foundation models, namely CLIP and SAM, we extract free pixel level annotations. We then propose a lightweight alignment module that projects the embedding of any arbitrary pretrained vision encoder to the text encoder space. By pretraining on COCO-Stuff, we obtain a generalist alignment module that can perform semantic segmentation across a wide range of object categories. Alternatively, our method can be deployed as a plug-and-play customized alignment module for any semantic segmentation dataset with zero annotations. We show SOTA results and open a new direction for zeroshot semantic segmentation by foundational models compositionality. As future work, we want to investigate other image encoders and continuous fine-tuning on new categories. Additionally, we want to explore how our alignment module can improve VLM models that are based on CLIP to further strengthen their object localization capabilities.